# On the dynamics of leapfrogging vortex rings

### v1

@@ 1. Introduction
Let us also mention that the strong localization $(ii)$ is obtained in a strip centered in $X_{i,0,r}$ for the sake of simplicity, one can easily adapt the proof to center the strip on $X_{i,r}(t)$ instead by actually centering the study on $b_{i,r}^{\varepsilon}(t)$.
@@
The plan of the paper is the following. In Section 2, we recall well-posedness results for (1.3) and provide a suitable expansion for the Biot-Savart law. In Section 3, we introduce useful notations and gather properties on the interactions of vortex rings. In Section 4, we obtain the weak localization property by new bootstrap argument and assuming the strong localization, in an unspecified region of space. In Section 5, we obtain information on the motion of the center of mass and obtain the weak localization Theorem 1.1-$(i)$. In Section 6, we conclude the proof of Theorem 1.1 by proving the strong localization, successively in the $r$ and $z$ direction.
@@ 2. Well-posedness and Biot-Savart law
2.1. Reduction
We denote by $(z,r,\theta)$ the usual cylindrical coordinates in ${\mathbb{R}}^{3}$ with basis vectors $(e_{z},e_{r},e_{\theta})$. A vector field $U:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{3}$ is called axisymmetric without swirl if of the form
$U(z,r,\theta)=U_{z}(z,r)e_{z}+U_{r}(z,r)e_{r},$
i.e. $U$ is independent of the toroidal variable $\theta$ (axisymmetric) and $U\cdot e_{\theta}=0$ (without swirl). This implies that the vorticity $\mathbf{\Omega}=\nabla\wedge U$ is a purely toroidal vector field:
$\mathbf{\Omega}=\Omega e_{\theta},\quad\text{with}\,\,\Omega=\partial_{z}U_{r}-\partial_{r}U_{z}.$
@@ 2.1. Reduction
$\omega=\frac{\Omega}{x_{r}},\qquad u=(U_{z},U_{r}).$
Note that the div-curl problem $\operatorname{div}U=0$ and $\nabla\wedge U=\mathbf{\Omega}$ reduces to
@@
Under suitable boundary and regularity conditions, the velocity field $u$ is hence uniquely determined by $\omega$, see Section 2.3 for details. On the half plane ${\mathbb{R}}_{+}^{2}$, the potential vorticity $\omega$ satisfies the transport type equation
@@ 2.2. Well-posedness results
Global existence and uniqueness of weak axisymmetric without swirl solutions of (1.1) has been proved in [50] and [46, Th 3.3], [1, 17] as well as references therein. In [14, Th. 5.4], global existence and uniqueness of the weak vorticity solution of (2.2) is proved when the initial velocity is axisymmetric without swirl, the initial vorticity satisfies $\Omega_{0}\in L^{1}\cup L^{\infty}({\mathbb{R}}_{+}^{2})$ and its support does not overlap the symmetry axis. More precisely, this solution satisfies
@@
In the solution constructed by [14, Th. 5.4], the velocity is reconstructed from the vorticity with the Biot-Savart Kernel: $U(t,x)=\int H(x-y)\wedge\mathbf{\Omega}(t,y)\,dy$ with $H(x)=-x/(4\pi|x|^{3})$ and with $\mathbf{\Omega}$ depending only on $(z,r)$ and purely toroidal. As used in [8, 12], the Biot-Savart law in this case admits an explicit representation:
[…]
Moreover, it is classical that $u$ is bounded (see (2.6) below) and log-lipschitz (by adapting, for instance, the planar case done in [41, App. 2.3]). Finally, it is established in [14, Corollary 5.6] that the solution is a weak solution of (2.2) in the sense of distributions (see (5.3) in [14]). We summarize these results on the half plane ${\mathbb{R}}_{+}^{2}$ in the next proposition.
@@
    The vorticity equation is satisfied in the sense of distribution, i.e. for all $\Phi\in C^{\infty}([0,T]\times{\mathbb{R}}_{+}^{2})$ and for any $t\in[0,T]$
@@
Let $\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ such that $x_{r}\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ be as in Proposition 2.1 above. Let $T>0$ and $(\omega,u)$ be the unique weak solution of (1.3) on $[0,T]$ provided by Proposition 2.1. Then the following holds true.
@@
    If $\omega_{0}$ has compact support in ${\mathbb{R}}_{+}^{2}$, denoting by $R_{0}=\sup\{|x|\,|\,x\in\operatorname{supp}(\omega_{0})\}$, there exists a compact set $K_{T}\subset{\mathbb{R}}_{+}^{2}$ depending only on $\|x_{r}\omega_{0}\|_{L^{\infty}}$, $\delta_{0},R_{0}$ and $T$ such that
    $\operatorname{supp}\omega(t,\cdot)\subset K_{T},\quad\forall t\in[0,T].$
@@
For initial vorticity as in (1.7) consisting of $N$ disjoint blobs, the existence of a unique Lagrangian flow allows one to decompose the unique weak solution as
@@
Moreover, Part (i) of Corollary 2.2 states that there exists a positive time such that a vorticity that is initially localized away from the symmetry axis does not reach the axis and remains localized. Concerning the localization in the radial direction, a quantitative estimate of this type is stated in [15, Section 2.3], it is shown that the radius of the support
@@
$R_{\varepsilon}(t)\leqslant C_{\varepsilon}(1+t)^{2},$

### v2

@@ 1.1. Main result
1.2. The limiting system
Several remarks concerning System (1.3) are in order. While it is Hamiltonian, it is not known whether solutions are global due to a potential collapse of vortex rings at $r=0$. We prove that there exists $T_{X}>0$ only depending on $(Y_{i,0})_{1\leqslant i\leqslant N}$ defined in (1.4) but not on $\varepsilon$ such that System (1.3) has a unique solution on the time interval $[0,T_{X}]$, see Section 5.2. We recall that all vortex rings are chosen to be of circulation $\gamma$ and hence it is convenient to consider the problem in a vertically moving frame in terms of $(Y_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ that are defined by the relation
(1.7) $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t).$
Further, we show that with that definition, the trajectories $t\mapsto Y_{i}^{\varepsilon}(t)$ converge to limiting trajectories $t\mapsto\widetilde{Y_{i}}(t)$ as $\varepsilon\to 0$ which satisfy
[…]
In this system, the relative self-induced motion and the interactions are readily seen to be of the same order. Note the negative sign of the second term (1.8) linked to the fluctuations in the relative self-induced vertical motion. The larger the radius $\widetilde{Y}_{i,r}$ the larger is the deceleration compared to the uniform vertical translation, see Appendix B for a detailed discussion of the limit dynamics. In the literature, such as [39, 20, 48], it is this dynamics that is considered as the motion of the rings in the appropriate scaling.
In this paper, we prefer to work with the trajectories $(X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ and $(Y_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ as it is much more natural in the computations. We prove in Section 5 that subsequently, one could replace the trajectories $t\mapsto X_{i}^{\varepsilon}(t)$ by
to match the notations of [39, 20, 48]. All the results, in particular Theorem 1.1 remain valid with that substitution.
On a related note, we point out that the limiting dynamics (1.8) differs qualitatively and quantitatively from (1.3). While, we can rule out collisions of points in (1.8) in finite time, see Appendix C.1, global existence for solutions to (1.3) remains open. For further differences in the behavior of solutions for $0<\varepsilon\ll 1$ and $\varepsilon=0$ including numerical evidence, see Appendix C.2. Moreover, we recover the same limit dynamics (1.8) as for the leapfrogging in [20] and for the Gross-Pitaevskii equation in [39] even though the dynamics for $\varepsilon>0$ in [39] appears to differ from the present one. Finally, we emphasize that neither the dynamics for $\varepsilon=0$ nor $\varepsilon>0$ considered here coincides with the ones studied in [11]. Specifically, the self-induced motion (not the relative one) is chosen to be of the same strength than the displacement due to interactions. In the respective limit dynamics (B.4) for $\varepsilon=0$, the displacement linked to the vertical self-induced motion does not depend on the specific radius $Y_{i,r}$. We refer to Appendix B for details.
1.3. Survey of previous results and outlook
The interest in the motion of vortex rings and their hydrodynamic stability dates back to the work of Lord Kelvin [40] and Helmholtz [33]. Rigorous derivations of the motion of a single vortex ring in an inviscid ideal fluid are provided in [29, 5]. The respective motion law to (1.1) for a single vortex ring was rigorously established for weakly viscous fluids in [7, 9, 31] where the recent paper [31] provides a very accurate description for large times.
More generally, assuming that the vorticity is concentrated around a curve, the inviscid motion is expected to be asymptotically governed by the binormal curvature flow equation, as was formally computed by Da Rios [54], see [37, 28] for partial results in this direction. The binormal curvature flow exhibits a rich panel of dynamics that have been intensively studied. In particular a series of papers by Banica and Vega is devoted to self-similar solutions or infinite polygonal lines solutions (we refer to [4] for a review of the results). Finally, Jerrard and Smets [38] introduced a new formulation of the binormal curvature flow equation allowing for weaker solutions.
The issue of the persistence in time of such vortex filaments by the Euler flow is the so-called vortex filament conjecture. While the conjecture remains open in its generality, results are available for specific geometric and dimension reduced settings. The rigorous asymptotic dynamics of straight filaments (the motion of Dirac masses of vorticity in the 2D Euler equations) was obtained in [44]. In the aforementioned case of annular vortex rings evolving according to (2.4), positive answers have been given with various assumptions. We refer to the work by Buttà, Cavallaro and Marchioro [8, 11] and references therein. This is in particular due to the fact that the problem allows for a dimension reduction leading to the 2D problem (1.2).
We emphasize that System (1.2) notably differs from the 2D Euler equations. Indeed, the dynamics of 2D Euler point-vortices do not include any self-induced motion of vortices. Even though Problem (1.2) is reduced to two dimensions, as in the study of helical filaments in [23] or point-vortices for the lake equations [35], the Biot-Savart law (the relation giving $u$ in terms of $\omega$) is much more complex compared to the planar case. Additional techniques are required and the results developed for the 2D Euler equations (see for instance [44]) cannot be applied. Roughly speaking, the motion of vortex rings is driven by the self-induced motion (1.1) and by their mutual interactions. Depending on the chosen asymptotic regime, the former maybe dominant, leading to a translation of the vortex rings along the symmetry axis [12, 8]. In case the latter dominates, the dynamics is reminiscent to the 2D dynamics of point vortices [10]. In particular, it was shown in [8] that the interactions of vortex rings of different radii (independent of $\varepsilon$) are bounded independently of $\varepsilon$ and can be neglected compared to their self-interaction (inducing the velocity given by (1.1)) which is of order $|\ln\varepsilon|$. On a related note, in the case of a single vortex ring, the result of Theorem 1.1 also improves those of [8].
As discussed, the leapfrogging phenomenon first described by Helmholtz [33] concerns the regime where both effect are balanced and hence requires mutual distances of order $1/\sqrt{|\ln\varepsilon|}$ to each other, and a detailed explanation of the suitable scaling regime is provided in Appendix B.2. The small mutual distance and strong interactions make the scenario hardly accessible by previously mentioned techniques.
To the best of our knowledge, the only rigorous results concerning the leapfrogging dynamics for the 3D Euler equations are [20, 11] and [39] for the 3D Gross-Pitaevskii equation. While [20, 39] consider the same scaling regime, the regime studied in [11] differs from the present one. The results from [11] cannot be adapted to our situation, as also pointed out by the authors in [11, page 791], see also Appendix B.2 for a comparison. In [20], the authors construct initial data $\omega_{0}^{\varepsilon}$ of a very specific class to obtain a smooth solution exhibiting leapfrogging vortex rings. We emphasize the fact that the conclusions of Theorem 1.1 hold for any initial data $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying relations (1.5), which ranges from smooth compactly supported blobs to vortex patches, even possibly multiply connected. Opposite to [20], our proof does not rely on sophisticated elliptic desingularization techniques.
While our result is limited to a fraction of a periodic motion but treats very general initial data, [20, 39] provide solutions exhibiting several crossings. The method in [39] for the Gross-Pitaevskii equation is based on Jacobian estimates not available for (2.1). In [11], several crossings are observed upon choosing large radii but in a different scaling regime and shorter time scales.
In view of these results, additional assumptions on the general class of initial data (1.5) considered here might be necessary. Indeed, it is conceivable that the possible occurrence of filamentation of the vortex ring in the vertical direction leads to instabilities preventing an improvement of the $\mathcal{O}(1)$ localization in the $z$ direction for general initial data. Namely, a tail in the wake of the ring may appear, a phenomenon that was described for instance in [49, 50] and recently proved in [17, 18, 46]. This fact may explain similar restrictions in [8, 12, 23, 35]. In [20], this filamentation is not present since the authors work with specific initial data and a stable vorticity profile is constructed. In [11], this filamentation does also not occur suggesting that the tail does not appear on the short time scale considered. Consequently, it appears to be of great relevance to investigate the appearance of vortex filamentation and related potentially destabilizing effects.
Remark 1.3.
We expect that our small time assumption in Theorem 1.1 is not sharp and can be optimized by revisiting the proof. However, it becomes clear from our proof that our method is not expected to reach a full rotation of two vortex rings around each other. The localization in the $x_{z}$ direction of order $\mathcal{O}(1)$ constitutes the main obstruction as the strips in which each ring is supported are not allowed to overlap. The possible occurrence of vortex filamentation discussed above appears to prevent a more precise localization.
Nevertheless, it might be possible to remove the small time assumption at least for the weak localization: this would require the weak localization to be obtained without assuming a priori strong localization, which seems not to be within reach of our current method. Indeed, also previous results for point vortices in inviscid $2D$ flows require a separation of vortex blobs [44]. This observation further motivates to study the stability of a vortex ring under collision with a tail of another vortex ring. However, this challenging mathematical problem is expected to require new tools.
1.4. Strategy of the proof
The 3D axisymmetric Euler equations without swirl allow for a dimension reduction in terms of a transport type equation for the vorticity coupled with an anelastic constraint for the velocity field (1.2). The self-induced motion of a vortex ring is closely linked to this anelastic constraint and also captured by the local induction approximation (LIA), see also [35, 23]. While this constitutes a main difference to the 2D setting, it is well known that the Biot-Savart law in the present setting shares some similarities with the 2D one even though it admits no explicit representation formula. Indeed, we revisit the well-posedness of (1.2) in Section 2 and provide a suitable expansion of the Biot-Savart law. This allows for a decomposition of the velocity field displaying the self-induced motion and mutual interactions of vortex rings (2.16). Based on this decomposition and a boot-strap hypothesis on the strong localization up to some time $T_{\varepsilon}>0$, we provide suitable estimates on the strength of the interactions of vortices. A common feature in previous papers [8, 35, 23] is that vorticity which is initially localized in balls of radius $\varepsilon$ is only known to be localized in balls of radius $|\ln\varepsilon|^{-\kappa}$ for some $40<\kappa<1/$ 4 for positive times. Given the mutual distance of $|\ln\varepsilon|^{-\frac{1}{2}}$ this turns out to be insufficient for our purpose and needs to be significantly improved by upgrading both weak and strong localization properties.
The next step consists in showing the weak localization property. Commonly and in the aforementioned results, this property is inferred for the $i$-th vortex by exploiting the conservation of $L^{p}$-norms of the potential vorticity, the total energy as well as suitable bounds on the local energy of the $i$-th vortex and higher order moments such as the moment of inertia. In particular, the use of the moment of inertia $I_{\varepsilon}(t)=\int|x-b_{\varepsilon}(t)|^{2}x_{r}\omega_{\varepsilon}(t)\mathrm{d}x$ causes a loss of precision in the localization for positive times due to a repeated application of Hölder-inequalities. To overcome this issue, we revisit the variational approach by Turkington [55] developed for a single vortex evolving according to the 2D Euler flow. Given a precise expansion of the local energy and the conservation of the mass of the $i$-th vortex, it allows one to precisely localize the vorticity in weak sense. Several key difficulties arise: the techniques need to be adapted to the present setting with an anelastic constraint and several vortices displaying singular interactions. As these interactions may potentially lead to large variations in the energy of the $i$-th vortex, a precise estimate on the respective energy as required by Turkington is difficult to obtain. We overcome these difficulties among others by a multi-layered boot-strap argument. Specifically, assuming additionally a bootstrap hypothesis on the weak localization up to some time $0<\widetilde{T_{\varepsilon}}\leqslant T_{\varepsilon}$, we are able to provide precise estimates on the center of vorticity and the energy of the $i$-th vortex. With this information at hand, we show a substantially more precise weak localization property by adapting Turkington’s argument. In particular, it follows $\widetilde{T_{\varepsilon}}=T_{\varepsilon}$. We believe the extension of Turkington’s approach to the setting of an anelastic constraint and strongly interacting vortices to be of independent interest.
A first consequence of this property is the derivation of the motion law up to time $T_{\varepsilon}$ in Section 5. Finally, we prove the desired strong localization property in Section 6. While no self-induced motion linked to the LIA occurs in the purely 2D setting, here it amounts to a fast vertical translation (1.1). This anisotropic behavior together with the possible appearance of vortex filamentation suggests that a strong localization on scales of order $\mathcal{O}(\sqrt{|\ln\varepsilon|}^{-1})$ for generic solutions can only be expected in the radial direction. Indeed, a small filament emanating from a vortex rings is expected to travel at speeds of order $\mathcal{O}(1)$ while the core of vortex rings travels at $\mathcal{O}(|\ln\varepsilon|)$ in $x_{z}$-direction. Concerning the strong localization in $x_{r}$-direction, we require a localization of the support of order $\mathcal{O}(|\ln\varepsilon|^{-\frac{1}{2}})$. In the literature, the proof of the strong localization is typically based on an iterative method first introduced by Marchioro and Pulvirenti in [44, 45] for $2D$ Euler flows. Up to today, the approach has reached an apparent complexity by being gradually upgraded and refined in a series of papers to various dimension-reduced settings and interacting vortices. Notably, in [8] and subsequently in [35, 23] the anisotropic character of the problem has been dealt with by introducing directional localization estimates. Most recently, a double iterative procedure has been introduced in [11] to show an improved strong localization result sufficient to deal with a leapfrogging type dynamics, though on a shorter time scale, see Appendix B for a comparison. Our aim is twofold. First, we adapt this double iterative procedure to prove the strong localization at the required precision up to time $T_{0}$ independent of $\varepsilon$, where we crucially rely on the improved weak localization estimates of Section 4. Second and as a service to the reader not familiar with the refined iteration procedure, we present the method in a self-contained and accessible manner.
1.5. Organization of the paper
@@
The plan of the paper is the following. In Section 2, we recall well-posedness results for (1.2) and provide a suitable expansion for the Biot-Savart law. In Section 3, we introduce useful notations and gather properties on the interactions of vortex rings. In Section 4, we obtain the weak localization property by new bootstrap argument and assuming the strong localization, in an unspecified region of space. In Section 5, we obtain information on the motion of the center of mass and obtain the weak localization Theorem 1.1-$(i)$. In Section 6, we conclude the proof of Theorem 1.1 by proving the strong localization, successively in the $r$ and $z$ direction. Appendix A provides two technical lemmas concerning rearrangement inequalities and a variant of Gronwall’s inequality. Appendices B-C discusses various asymptotic dynamics and provides a comparison to the regimes considered in [11, 20, 39].
@@ 2. Well-posedness and Biot-Savart law
2.1. 2D reduction
The evolution of an incompressible and inviscid three-dimensional fluid is governed by the standard 3D Euler equations
where $\mathbf{U}$ is uniquely determined by $\mathbf{\Omega}$ for all $t\in[0,T]$ by means of the Biot-Savart law, i.e. the second identity in (2.2). In the present paper, we restrict our attention to solutions that are axisymmetric vector fields without swirl. Thus, it is conveniently studied in cylindrical coordinates in ${\mathbb{R}}^{3}$ which we denote by $(z,r,\theta)$ with basis vectors $(e_{z},e_{r},e_{\theta})$. A vector field $\mathbf{U}:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{3}$ is called axisymmetric without swirl if of the form
$\mathbf{U}(z,r,\theta)=U_{z}(z,r)e_{z}+U_{r}(z,r)e_{r},$
$\mathbf{\Omega}=\Omega e_{\theta},\quad\text{with}\,\,\Omega=\partial_{z}U_{r}-\partial_{r}U_{z}\in{\mathbb{R}}.$
@@
$\omega=\frac{\Omega}{x_{r}}\in{\mathbb{R}},\qquad u=(U_{z},U_{r})\in{\mathbb{R}}^{2}.$
Note that the div-curl problem $\operatorname{div}\mathbf{U}=0$ and $\nabla\wedge\mathbf{U}=\mathbf{\Omega}$ reduces to
@@
Under suitable boundary and regularity conditions, the two-dimensional velocity field $u$ is hence uniquely determined by $\omega$, see Section 2.3 for details. On the half plane ${\mathbb{R}}_{+}^{2}$, the potential vorticity $\omega$ satisfies the transport type equation
@@ 2.2. Well-posedness results
Global existence and uniqueness of weak axisymmetric without swirl solutions of (2.1) has been proved in [56] and [53, Theorem 3.3], [1, 19] as well as references therein. In [14, Theorem 5.4], global existence and uniqueness of the weak vorticity solution of (2.4) is proved when the initial velocity is axisymmetric without swirl, the initial vorticity satisfies $\Omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ and its support does not overlap the symmetry axis. More precisely, this solution satisfies
@@
In the solution constructed by [14, Theorem 5.4], the velocity is reconstructed from the vorticity with the Biot-Savart Kernel: $U(t,x)=\int H(x-y)\wedge\mathbf{\Omega}(t,y)\,\mathrm{d}y$ with $H(x)=-x/(4\pi|x|^{3})$ for $x,y\in{\mathbb{R}}^{3}$ and with $\mathbf{\Omega}$ depending only on $(z,r)$ and purely toroidal. Writting this 3D Biot-Savart law in cylindrical coordinates, we have an explicit representation [^1] We may note that a $\cos\theta$ is missing in [8, 12].:
(2.7) $\begin{split}u_{z}(t,x)&=-\frac{1}{4\pi}\int\,\mathrm{d}y_{z}\int_{0}^{\infty}y_{r}^{2}\,\mathrm{d}y_{r}\int_{-\pi}^{\pi}\,\mathrm{d}\theta\frac{\omega(t,y)(x_{r}\cos\theta-y_{r})}{[|x-y|^{2}+2x_{r}y_{r}(1-\cos\theta)]^{3/2}},\\ u_{r}(t,x)&=\frac{1}{4\pi}\int\,\mathrm{d}y_{z}\int_{0}^{\infty}y_{r}^{2}\,\mathrm{d}y_{r}\int_{-\pi}^{\pi}\,\mathrm{d}\theta\frac{\omega(t,y)(x_{z}-y_{z})\cos\theta}{[|x-y|^{2}+2x_{r}y_{r}(1-\cos\theta)]^{3/2}},\end{split}$
where $x=(x_{z},x_{r})$ and $y=(y_{z},y_{r})$ are vectors in +${\mathbb{R}}^{2}_{+}$. Moreover, it is classical that $u$ is bounded (see (2.9) below) and log-lipschitz (by adapting, for instance, the planar case done in [45, Appendix 2.3]). Finally, it is established in [14, Corollary 5.6] that the solution is a weak solution of (2.4) in the sense of distributions (see (5.3) in [14]). We summarize these results for
(2.8) $\begin{cases}\partial_{t}\,\omega+u\cdot\nabla\omega=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}(x_{r}u)=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u)=x_{r}\omega&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega(0,\cdot)=\omega_{0}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
on the half plane ${\mathbb{R}}_{+}^{2}$ in the next proposition.
@@
    The vorticity equation is satisfied in the sense of distribution, i.e. for all $\Phi\in C^{\infty}_{c}([0,T]\times{\mathbb{R}}_{+}^{2})$ and for any $t\in[0,T]$
@@
Let $\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ such that $x_{r}\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ and $\delta_{0}:=\inf_{x\in\operatorname{supp}(\omega_{0})}x_{r}>0$. Let $T>0$ and $(\omega,u)$ be the unique weak solution of (2.8) on $[0,T]$ provided by Proposition 2.1. Then the following holds true.
@@
    If $\omega_{0}$ has compact support in ${\mathbb{R}}_{+}^{2}$, denoting by $R_{0}=\sup\{|x|\,|\,x\in\operatorname{supp}(\omega_{0})\}$, there exists a compact set $\mathcal{K}_{T}\subset{\mathbb{R}}_{+}^{2}$ depending only on $\|x_{r}\omega_{0}\|_{L^{\infty}}$, $\delta_{0},R_{0}$ and $T$ such that
    $\operatorname{supp}\omega(t,\cdot)\subset\mathcal{K}_{T},\quad\forall t\in[0,T].$
@@
Finally, one obtains the scaled system (1.2) from (2.8) by considering the logarithmic time scale $s=|\ln\varepsilon|t\in[0,T]$ and $\omega^{\varepsilon}(s,x)=\omega\left(\frac{s}{|\ln\varepsilon|},x\right)$.
Further, for initial vorticity as in (1.5) consisting of $N$ disjoint blobs, the existence of a unique Lagrangian flow allows one to decompose the unique weak solution as
@@
Moreover, Part (i) of Corollary 2.2 states that there exists a positive time such that a vorticity that is initially localized away from the symmetry axis does not reach the axis and remains localized. Concerning the localization in the radial direction, a quantitative estimate of this type is stated in [17, Section 2.3], it is shown that the radius of the support
@@
$R_{\varepsilon}(t)\leqslant C_{\varepsilon}\Big{(}1+\frac{t}{|\ln\varepsilon|}\Big{)}^{2},$